# Will anyone review this paper? Screening, sorting, and the feedback cycles that imperil peer review

### v2

@@ III.1 A single journal
A central result of AO is that more accurate reviewing results in fewer authors submitting their paper, while less accurate reviewing leads to more submissions (Fig. 2C). The intuition is that more accurate peer review tightens the relationship between the author’s acceptance probability and their type, which strengthens screening. Thus more accurate reviewing makes both authors and reviewers better off because it decreases $L$. The effect on readers’ welfare is ambiguous.
@@
If increasing the review load also decreases the accuracy of peer review, then author screening and journal sorting participate in a feedback loop [56]. It is easy to see how an increase in the review load might cause peer review to become less accurate. Suppose that reviewers differ in their review accuracy and that editors preferentially invite more accurate reviewers. As the review load increases, editors must either ask their favored reviewers to do more work or they must reach further into the pool of reviewers. In either case, the accuracy of peer reviews declines. Thus, review noise $\sigma_{Y}$ and review load $L$ reinforce each other: more accurate peer review compels authors to screen more selectively, and tighter screening results in fewer submitted manuscripts and more accurate peer review. In the other direction, noisier peer reviewing results in looser screening, which in turn creates more work for reviewers, leading to even noisier reviewing. (A formal mathematical statement of a model with this feedback loop appears in the appendix.)
@@
The feedback between sorting and screening affects how the equilibrium changes when the environment is perturbed. For example, in recent decades the career rewards from publishing in top-tier journals have increased dramatically [53, 32]. It is easy to show that when publishing in top journals brings greater rewards—that is, $v/c$ increases—more authors throw their hat in the ring and the review load increases (Fig. 2D). The intuition is obvious: when publication is worth more, authors are willing to take a chance on a lower probability of success. [^4] The welfare effects of an increase in $v/c$ are mixed. Clearly, the increase in the review load imposes an additional burden on reviewers. The welfare consequences of an increase in $v/c$ to authors are ambiguous, however, because even though more authors are paying the cost of submission, submission is relatively less costly, and thus the net effect could be positive or negative. The effect of an increase in $v/c$ on the quality of published articles is also ambiguous.
@@
The increase in the review load $L$ caused by an increase in $v/c$ is exacerbated by the feedback loop between $L$ and the review error $\sigma_{Y}$. Fig. 4A,B illustrates. (Fig. 4C will be considered later.) In each panel of Fig. 4, the red, blue, and black curves show how the review load $L$ responds to a change in the review noise $\sigma_{Y}$, for different values of $v/c$. The dashed green curve shows how $\sigma_{Y}$ might respond to the review load $L$. The equilibrium for a particular $v/c$ is found at the intersection of the green curve with the respective red, blue, or black curves. Comparing Fig. 4A to Fig. 4B shows that when review accuracy declines with $L$, an increase in $v/c$ increases the review load more than it would if reviewer accuracy was independent of $L$. In other words, not only does an increase in $v/c$ compel more authors to submit their paper, the resulting pressure on the review pool decreases the accuracy of review, resulting in yet more authors submitting their manuscripts. [^5] In principle, the feedback between sorting and screening could cause the equilibrium to fully unravel, only stopping when every author submits every manuscript. In Fig. 4, such an unraveling would occur if the dashed green curve for $\sigma(L)$ intersects the red, blue, or black curves from below. We don’t think such a scenario is likely under reasonable model settings, however, and experience suggests that the peer-review process has not yet completely unraveled.
@@
The feedback loop between $\sigma_{Y}$ and $L$ also prolongs transient perturbations to the equilibrium caused by one-time shocks. Although a formal dynamical model is outside the present scope, the intuition for the effect is easily seen. Suppose there is a one-time shock that increases the review load to a value $L^{\prime}$ above its equilibrium level. If the review noise depends on $L$, then the review noise will increase in response to the surge in submissions to a value $\sigma^{\prime}_{Y}$. After the shock is over, the submission load does not immediately return to its pre-shock value; instead, it adjusts to the value $L^{\prime\prime}<L^{\prime}$ induced by review noise $\sigma^{\prime}_{Y}$. The review noise then adjusts to the value $\sigma^{\prime\prime}_{Y}$ induced by $L^{\prime\prime}$, and so on. Eventually, $L$ and $\sigma_{Y}$ will return to their pre-shock equilibrium, but these transients will take longer to relax if the review noise temporarily increases in response to the one-time surge in submissions. This suggests one plausible explanation for why a reported change in reviewer behavior following the COVID-19 pandemic [23] may be surprisingly persistent.
@@ III.2 Several competing journals
Here, we consider how the shift from a few large journals to several smaller ones has impacted scientists’ welfare. We focus on “horizontal” proliferation of journals within a prestige tier. “Vertical” proliferation, in which journals more densely occupy niches along a prestige spectrum, is outside the present scope (but see ref. [41]). We also abstract away from ways in which journals within a prestige tier may specialize to particular subsets of scholarship.
@@
To develop notation, write the conditional probability that author $q$’s manuscript is accepted on the $j$-th attempt when facing review-score threshold $y$ as
$a_{j}(q;y)=\mathsf{Pr}\!\left\{Y_{j}\geq y|F_{X}(X)=q,\;Y_{1}<y,\dots,Y_{j-1}<y\right\}\\ j=1,\dots,J\,.$
While $a_{j}(q;y)$ does not depend directly on $J$, the equilibrium review cut-off $\hat{y}_{J}$ will depend on $J$, and thus $a_{j}(q;\hat{y}_{J})$ will as well. Write the marginal probability that author $q$’s manuscript is accepted on the $j$-th attempt as
$b_{j}(q;y)=\mathsf{Pr}\!\left\{Y_{1}<y,\ldots,Y_{j-1}<y,Y_{j}\geq y|F_{X}(X)=q\right\}\\ j=1,\dots,J\,.$
where $b_{j}(q;y)=0$ if it is not worthwhile for author $q$ to submit their manuscript a $j$-th time. Write the probability that author $q$’s manuscript is eventually accepted as $m_{J}(q;y)=\sum_{j=1}^{J}b_{j}(q;y)$. The total volume of published manuscripts at each period is then $\int_{0}^{1}\,m_{J}(q;y)\,dq$. An equilibrium consists of a review-score cutoff $\hat{y}_{J}$ at which the journals exactly fill their aggregate capacity and a set of marginal authors $\hat{q}_{1}\leq\hat{q}_{2}\leq\ldots\leq\hat{q}_{J}$ for which author $\hat{q}_{j}$ is indifferent about submitting their manuscript for the $j$-th time. The marginal authors satisfy the author-rationality conditions
$a_{j}(\hat{q}_{j},\hat{y}_{J})=c/v$ (3)
while the capacity-filling condition writes as
$\int_{0}^{1}\,m_{J}(q;\hat{y}_{J})\,dq=k.$ (4)
To calculate the review load, write the average number of times that author $q$ submits their manuscript as $\mu_{J}(q)$, given by
$\mu_{J}(q)=\text{1}_{q\geq\hat{q}_{1}}+\text{1}_{q\geq\hat{q}_{2}}\times(1-a_{1}(q,\hat{y}_{J}))\\ +\cdots+\text{1}_{q\geq\hat{q}_{J}}\times\prod_{j=1}^{J-1}\bigl{(}1-a_{j}(q,\hat{y}_{J})\bigr{)}.$
where $\text{1}_{q\geq\hat{q}_{i}}$ is an indicator function that equals 1 if $q\geq\hat{q}_{i}$ and equals 0 otherwise. (Note that $\mu_{J}(q)$ also gives the aggregate density of incoming submissions across all $J$ journals from type-$q$ authors at each period.) The total volume of submissions, and hence the review load, at each period is then
@@
Welfare measures extend naturally. Reader welfare is given by the average quality of published papers standardized by the quality of the top $k$ papers (a formal expression appears in the appendix), and reviewer welfare is proportional to (the opposite of) the review load $L_{J}$. Author $q$’s payoff equals $v\,m(q)-c\,\mu(q)$; integrating over authors again gives $\int_{0}^{1}\,[v\,m(q)-c\,\mu(q)]\,dq=v\,k-c\,L$; thus standardized author welfare is still 1cL/vk1-c\,L/v\,k.
The feedback loop between the quality of sorting and the strength of screening carries forward to this multi-journal model, and the underlying logic remains the same. We focus here on new phenomena caused by the presence of several journals, and study the model with a numerical example. This example suggests that increasing the number of journals results in journals publishing better papers while also increasing the review load and the volume of rejected manuscripts. Fig. 2E shows how this happens. More journals create more opportunities for authors to have their work considered afresh. More such opportunities increase the volume of recirculated manuscripts, which in turn forces journals to be more selective. Increased selectivity results in higher rejection rates which increases the volume of recirculated manuscripts, and so on. Thus, giving authors more bites at the apple strengthens screening in the sense that fewer authors find it worthwhile to submit their paper in the first place ($\hat{q}_{1}$ increases with $J$), but because those authors can submit their work several times, the sorting burden grows, and the overall volume of submitted (and rejected) manuscripts rises (Fig. 5).
@@
Moreover, increasing the number of journals has a bigger effect on the review load when reviews are noisy (Fig. 5). The intuition here is that authors learn less about their manuscript’s quality from journal decisions when reviews are noisy than they do when reviews are accurate. Thus, when reviews are noisy, more authors rationally continue to submit the same manuscript despite previous rejections. Combined with the feedback loop between sorting and screening discussed earlier, this identifies yet another feedback cycle by which the burden on peer reviewers grows: as the number of journals proliferates, authors have more opportunities to have their work considered afresh; more such opportunities increase the load on the review community; an increased review load results in less accurate reviews; declining review accuracy diminishes authors’ ability to learn, leading more authors to recycle previously rejected manuscripts, and so on. Fig. 4C illustrates how an increase in the number of journals amplifies the feedback loop between sorting and screening.
To see the effect of journal proliferation starkly, consider the extreme case in which authors know their manuscript’s quality exactly ($\sigma_{X}=0$; a similar analysis of this extreme case appears in ref. [56]). In this case, an author’s acceptance probability at any given attempt is independent of the number of times the manuscript has already been rejected ($a_{j}(q;y)=a(q;y)$ for all $j$). Thus there is a single marginal author, and an author’s eventual probability of acceptance writes neatly as $m_{J}(q;y)=1-(1-a(q;y))^{J}$. In the limit as $J$ grows large, any paper that worth submitting gets published eventually ($\lim_{J\rightarrow\infty}m_{J}(q;y)=1$). Writing the limiting marginal author as $\hat{q}_{\infty}$, the capacity-filling condition becomes
@@
In other words, when perfectly informed authors have an unlimited number of attempts to submit their work, screening is perfect: only the top $k$ authors ever submit their manuscript, and they keep submitting until their manuscript is published (Fig. 2F). [^6] This might also explain why elite journals are somewhat willing to give top authors an opportunity to appeal for fresh reviews of rejected manuscripts. At least we’ve heard that they are. The entire function of peer review in this case is to drive author 1k1-k’s probability of acceptance down to $c/v$ so that no other authors find it worthwhile to submit. The sorting function of peer review is just necessary waste, because every submitted manuscript is published eventually. Further, the equilibrium rejection rate increases with both $v/c$ and $\sigma_{Y}$. If $v/c$ rises, the acceptance threshold $\hat{y}_{\infty}$ must be pushed even higher to discourage non-submitting authors, leading to more superfluous rejection and resubmission of top papers. If $\sigma_{Y}$ increases, the acceptance probabilities of all authors above the marginal author decrease, forcing them to submit their paper more often before it is finally accepted.
@@ III.3 Desk rejection
As screening and sorting weaken, eventually a journal is bound to receive more manuscripts than it can suitably review. Of course, journals are not obligated to send every manuscript out for review. Instead, journals can and do counter a surge in submissions by desk-rejecting some manuscripts, thus preserving their available review labor for the most promising submissions [43, 30]. Yet anticipating the net effect of desk-rejection is complicated, because, as we have observed, authors choose whether or not to submit their manuscript in anticipation of the sorting process to follow. Presumably, even an informed journal editor will make less accurate decisions without peer reviews than with them, and so it is unclear on its face how desk rejection will affect screening, and how this will in turn affect the quality of a journal’s published articles. Here, we add desk-rejection to the one-journal model to formalize this idea and briefly explore its potency. A formal model of desk-rejection with competing peer journals is beyond the scope of the present article, but we informally discuss how competition impinges on desk rejection later.
@@ III.3.1 A single journal
For any possible desk-rejection threshold $d$, there is a corresponding marginal author $\hat{q}(d)$ and review score threshold $\hat{y}(d)$ that satisfy the author-rationality and capacity-filling conditions. The journal’s utility $u(d)$ is the average quality of accepted manuscripts, which writes as
@@
With desk rejection in play, use $S=1-\hat{q}$ to denote the volume of submitted manuscripts, and continue to use $L$ to denote the volume of manuscripts sent out for review. Author welfare is now given by 1cS/vk1-c\,S/v\,k, while reviewer welfare remains proportional to (the opposite of) $L$.

### v3

@@ III.1 A single journal
A central result of AO is that more authors submit their paper when reviewing is less accurate, and vice versa (Fig. 2C). The intuition is that less accurate peer review introduces more stochasticity into the journal’s sorting. This loosens the relationship between the author’s acceptance probability and their type, thus encouraging more authors to submit their manuscripts and relaxing screening. As a result, less accurate reviewing makes both authors and reviewers worse off by increasing the review load $L$. The effect on readers’ welfare is ambiguous.
@@
If increasing the review load also decreases the accuracy of peer review, then author screening and journal sorting participate in a feedback loop [41]. It is easy to see how an increase in the review load might cause peer review to become less accurate. Suppose that reviewers differ in their review accuracy and that editors preferentially invite more accurate reviewers. As the review load increases, editors must either ask their favored reviewers to do more work or they must reach further into the pool of reviewers. In either case, peer reviews become less accurate, making the journal’s sorting less precise. Thus, review noise $\sigma_{Y}$ and review load $L$ reinforce each other: more accurate journal sorting compels authors to screen more selectively, and tighter screening results in fewer submitted manuscripts and more accurate peer review. In the other direction, less accurate journal sorting results in looser screening, which in turn creates more work for reviewers, leading to even noisier reviewing. (A formal mathematical statement of a model with this feedback loop appears in the appendix.)
@@
The feedback between screening and sorting affects how the equilibrium changes when the environment is perturbed. For example, in recent decades the career rewards from publishing in top-tier journals have increased dramatically [46, 47]. It is easy to show that when publishing in top journals brings greater rewards—that is, $v/c$ increases—more authors throw their hat in the ring and the review load increases (Fig. 2D). The intuition is obvious: when publication is worth more, authors are willing to take a chance on a lower probability of success. [^4] The welfare effects of an increase in $v/c$ are mixed. Clearly, the increase in the review load imposes an additional burden on reviewers. The welfare consequences of an increase in $v/c$ to authors are ambiguous, however, because even though more authors are paying the cost of submission, submission is relatively less costly, and thus the net effect could be positive or negative. The effect of an increase in $v/c$ on the quality of published articles is also ambiguous.
@@
The increase in the review load caused by an increase in $v/c$ is exacerbated by the feedback loop between screening and sorting. Fig. 4A,B illustrates by showing how the review load $L$ (determined by the strength of screening) and the review noise $\sigma_{Y}$ (which determines the accuracy of sorting) shape each other. (Fig. 4C will be considered later.) In each panel of Fig. 4, the red, blue, and black curves show how the review load $L$ responds to a change in the review noise $\sigma_{Y}$, for different values of $v/c$. The dashed green curve shows how $\sigma_{Y}$ might respond to $L$. The equilibrium for a particular value of $v/c$ is found at the intersection of the green curve with the respective red, blue, or black curves. When review accuracy declines with increasing $L$ (Fig. 4B), an increase in $v/c$ increases the review load more than it would if reviewer accuracy was independent of $L$ (Fig. 4A). In other words, not only does an increase in $v/c$ compel more authors to submit their paper, the resulting pressure on the review pool decreases the accuracy of review, resulting in yet more authors submitting their manuscripts. [^5] In principle, the feedback between sorting and screening could cause the equilibrium to fully unravel, only stopping when every author submits every manuscript. In Fig. 4, such an unraveling would occur if the dashed green curve for $\sigma(L)$ intersects the red, blue, or black curves from below. We don’t think such a scenario is likely under reasonable model settings, however, and experience suggests that the peer-review process has not yet completely unraveled.
@@
The feedback loop between screening and sorting also prolongs transient perturbations to the equilibrium caused by one-time shocks. Although a formal dynamical model is outside the present scope, the intuition for the effect is easily seen. Suppose there is a one-time shock that relaxes screening and increases the review load to a value $L^{\prime}$ above its equilibrium level. If the review noise depends on $L$, then the review noise will increase in response to the surge in submissions to a value $\sigma^{\prime}_{Y}$. After the shock is over, the submission load does not immediately return to its pre-shock value; instead, it adjusts to the value $L^{\prime\prime}<L^{\prime}$ induced by review noise $\sigma^{\prime}_{Y}$. The review noise then adjusts to the value $\sigma^{\prime\prime}_{Y}$ induced by $L^{\prime\prime}$, and so on. Eventually, $L$ and $\sigma_{Y}$ will return to their pre-shock equilibrium, but these transients will take longer to relax if the review noise temporarily increases in response to the one-time surge in submissions. This suggests one plausible explanation for why a reported change in reviewer behavior following the COVID-19 pandemic [14] may be surprisingly persistent.
@@ III.2 Several competing journals
Here, we consider how the shift from a few large journals to several smaller ones has impacted scientists’ welfare. We focus on “horizontal” proliferation of journals within a prestige tier. “Vertical” proliferation, in which journals more densely occupy niches along a prestige spectrum, is more difficult to handle because it introduces greater richness into the author’s possible actions and consequently lies outside the present scope (but see ref. [44]). We also abstract away from ways in which journals within a prestige tier may specialize to particular subsets of scholarship.
@@
To develop notation, write the conditional probability that author $q$’s manuscript is accepted on the $j$-th attempt when facing review-score threshold $y$ as {multline*} a_j(q;y) = $\mathsf{Pr}\!\left\{Y_{j}\geq y|F_{X}(X)=q,\;Y_{1}<y,\dots,Y_{j-1}<y\right\}$
j=1,…,J . While $a_{j}(q;y)$ does not depend directly on $J$, the equilibrium review cut-off $\hat{y}_{J}$ will depend on $J$, and thus $a_{j}(q;\hat{y}_{J})$ will as well. Write the marginal probability that author $q$’s manuscript is accepted on the $j$-th attempt as {multline*} b_j(q;y) = $\mathsf{Pr}\!\left\{Y_{1}<y,\ldots,Y_{j-1}<y,Y_{j}\geq y|F_{X}(X)=q\right\}$
j=1,…,J . where $b_{j}(q;y)=0$ if it is not worthwhile for author $q$ to submit their manuscript a $j$-th time. Write the probability that author $q$’s manuscript is eventually accepted as $m_{J}(q;y)=\sum_{j=1}^{J}b_{j}(q;y)$. The total volume of published manuscripts at each period is then $\int_{0}^{1}\,m_{J}(q;y)\,dq$. An equilibrium consists of a review-score cutoff $\hat{y}_{J}$ at which the journals exactly fill their aggregate capacity and a set of marginal authors $\hat{q}_{1}\leq\hat{q}_{2}\leq\ldots\leq\hat{q}_{J}$ for which author $\hat{q}_{j}$ is indifferent about submitting their manuscript for the $j$-th time. The marginal authors satisfy the author-rationality conditions
[…]
while the capacity-filling condition is given by
$\tagCFJ\int_{0}^{1}\,m_{J}(q;\hat{y}_{J})\,dq=k.\tag${CF-J} (4)
To calculate the review load, write the average number of times that author $q$ submits their manuscript as $\mu_{J}(q)$, given by {multline*} μ_J(q) = 1_q ≥^q_1 + 1_q ≥^q_2 ×(1 - a_1(q,^y_J))
+ ⋯+ 1_q ≥^q_J ×∏_j=1^J-1(1 - a_j(q,^y_J)). where $\text{1}_{q\geq\hat{q}_{i}}$ is an indicator function that equals 1 if $q\geq\hat{q}_{i}$ and equals 0 otherwise. (Note that $\mu_{J}(q)$ also gives the aggregate density of incoming submissions across all $J$ journals from type-$q$ authors at each period.) The total volume of submissions, and hence the review load, at each period is then
@@
Welfare measures extend naturally. Reader welfare is given by the average quality of published papers standardized by the quality of the top $k$ papers (a formal expression appears in the appendix). The burden on reviewers is proportional to the review load $L_{J}$. Author $q$’s payoff equals $v\,m(q)-c\,\mu(q)$; integrating over authors again gives $\int_{0}^{1}\,[v\,m(q)-c\,\mu(q)]\,dq=v\,k-c\,L$; thus standardized author welfare is still 1cL/vk1-c\,L/v\,k.
The feedback loop between screening and sorting carries forward to this multi-journal model, and the underlying logic remains the same. We focus here on new phenomena caused by the presence of several journals and study a numerical example. This example suggests that increasing the number of journals results in journals publishing better papers while also increasing the review load and the volume of rejected manuscripts. Fig. 2E shows how this happens. More journals create more opportunities for authors to have their work considered afresh. More such opportunities increase the volume of recirculated manuscripts, which in turn forces journals to be more selective. Increased selectivity results in higher rejection rates which increases the volume of recirculated manuscripts, and so on. Thus, giving authors more bites at the apple strengthens screening in the sense that fewer authors find it worthwhile to submit their paper in the first place ($\hat{q}_{1}$ increases with $J$), but because those authors can submit their work several times, the sorting burden grows, and the overall volume of submitted (and rejected) manuscripts rises (Fig. 5).
@@
Moreover, increasing the number of journals has a bigger effect on the review load when reviews are noisy (Fig. 5). The intuition here is that authors learn less about their manuscript’s quality from journal decisions when reviews are noisy than they do when reviews are accurate. Thus, when reviews are noisy, more authors rationally continue to submit the same manuscript despite previous rejections. Combined with the feedback loop between screening and sorting, this identifies yet another feedback cycle by which the burden on peer reviewers grows: as the number of journals proliferates, authors have more opportunities to have their work considered afresh; more such opportunities increase the load on the review community; an increased review load results in less accurate reviews; declining review accuracy diminishes authors’ ability to learn, leading more authors to recycle previously rejected manuscripts, and so on. Fig. 4C illustrates how an increase in the number of journals amplifies the feedback loop between screening and sorting.
To see the effect of journal proliferation starkly, consider the extreme case in which authors know their manuscript’s quality exactly ($\sigma_{X}=0$; a similar analysis of this extreme case appears in ref. [41]). In this case, an author’s acceptance probability at any given attempt is independent of the number of times the manuscript has already been rejected ($a_{j}(q;y)=a(q;y)$ for all $j$). Thus there is a single marginal author, and an author’s eventual probability of acceptance is $m_{J}(q;y)=1-(1-a(q;y))^{J}$. In the limit as $J$ grows large, any paper that worth submitting gets published eventually ($\lim_{J\rightarrow\infty}m_{J}(q;y)=1$). Writing the limiting marginal author as $\hat{q}_{\infty}$, the capacity-filling condition becomes
@@
In other words, when perfectly informed authors have an unlimited number of attempts to submit their work, screening is perfect: only the top $k$ authors ever submit their manuscript, and they keep submitting until their manuscript is published (Fig. 2F). [^6] This might also explain why elite journals are somewhat willing to give top authors an opportunity to appeal for fresh reviews of rejected manuscripts. At least we’ve heard that they are. The entire function of peer review in this case is to drive the marginal author’s probability of acceptance down to $c/v$ so that no other authors find it worthwhile to submit. The sorting function of peer review is just necessary waste, because every submitted manuscript is published eventually. Further, the equilibrium rejection rate increases with both $v/c$ and $\sigma_{Y}$. If $v/c$ rises, the acceptance threshold $\hat{y}_{\infty}$ must be pushed even higher to discourage non-submitting authors, leading to more superfluous rejection and resubmission of top papers. If $\sigma_{Y}$ increases, the acceptance probabilities of all authors above the marginal author decrease, forcing them to submit their paper more often before it is finally accepted.
@@ III.3 Desk rejection
As screening and sorting weaken, eventually a journal is bound to receive more manuscripts than it can review. Of course, journals are not obligated to send every manuscript out for review. Instead, journals can and do counter a surge in submissions by desk-rejecting some manuscripts, thus preserving their available review labor for the most promising submissions [49, 50]. Yet anticipating the net effect of desk-rejection is complicated, because, as we have observed, authors choose whether or not to submit their manuscript in anticipation of the sorting process to follow. Presumably, even an informed journal editor will make less accurate decisions without peer reviews than with them, and so it is unclear on its face how desk rejection will affect screening, and how this will in turn affect the quality of a journal’s published articles. Here, we add desk-rejection to the one-journal model to formalize this idea and briefly explore its potency. A formal model of desk-rejection with competing peer journals is beyond the scope of the present article, but we informally discuss how competition impinges on desk rejection later.
@@ III.3.1 A single journal
For any possible desk-rejection threshold $d$, there is a corresponding marginal author $\hat{q}(d)$ and review score threshold $\hat{y}(d)$ that satisfy the author-rationality and capacity-filling conditions. The journal’s utility $u(d)$ is the average quality of accepted manuscripts, which is given by the expression
@@
[…]
With desk rejection in play, use $S=1-\hat{q}$ to denote the volume of submitted manuscripts, and continue to use $L$ to denote the volume of manuscripts sent out for review. Author welfare is now given by 1cS/vk1-c\,S/v\,k, while the burden on reviewers continues to scale with $L$.